Cardinal Inc. will get a revenue of $8000 per quarter for 10 years at 12% per year compounded monthly. What is the future worth at the end of 10 years?
Answer #1

Quarterly payment (A) = 8000

r = 12% or 0.12

m = 4 * 10 years = 40

n = 12

t = 10 years

number of periodic payments in the compounding period (p) = 4 / 12 = 0.3333

Future Value = A* p {[(1 + r/n)(nt) - 1] / (r/n)}

                     = 8,000 * 0.3333{[(1 + 0.12 / 12)12*10 - 1] / (0.12 / 12)}

                     = 8,000 * 0.3333{[(3.3004 - 1] / (0.01)}

                     = 8,000 * 0.3333(2.3004 / 0.01)

                     = $613,378.66

Thus, the future worth at the end of 10 year is $613,378.66